ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

11. Mathematical Physics

1991 MS Classification: 57M25, 57M30, 57Q45, 81 Monastyrsky, M., Inst. of Theoreti cal and experimental Physics, Moscow, Russia Correlation functions on Riemann surfaces and statistics of Knots In this talk we discuss the connections between two dimensional conformal field theories, random walks on hyperbolic Riemann surfaces and knot theory. For the wide class of CFT with monodromy group being the discrete subgroup of SL(2, R) we determinate the four point correlation functions. This construction based on some relations between monodromy representation of braid groups and invariants of knots. The talk based on the joint works with S. Nechaev (IPN, ORSAY, France).